# Fiducial Inference for Random-Effects Calibration Models: Advancing Reliable Quantification in Environmental Analytical Chemistry

### v2

@@ 4.1.3 Updated Fiducial Quantities for $\alpha_{i}$
The method for constructing equations that lead to fiducial quantities for $\alpha_{i}$ is similar to that for $\beta_{i}$ (for $i=1,...,q$), as discussed in the previous subsection. In the following, we construct a total of $q$ equations, one for each $\alpha_{i}$, by summing equation (4) over all replications at different non-zero concentrations for each lab, $\sum\limits_{j=2}^{r}\sum\limits_{k=1}^{N_{ij}}\frac{y_{ijk}}{x_{j}}=\alpha_{i}N_{ij}\sum\limits_{j=2}^{r}\frac{1}{x_{j}}+\beta_{i}\sum\limits_{j=2}^{r}\sum\limits_{k=1}^{N_{ij}}\exp(\sigma_{\eta}z_{ijk}^{\eta})+\sigma_{\epsilon}\sum\limits_{j=2}^{r}\sum\limits_{k=1}^{N_{ij}}\frac{z_{ijk}^{\epsilon}}{x_{j}},\ i=1,...,q.$ Again, these are linear in each $\alpha_{i}$, $i=1,...,q$. We now replace $\beta_{i}$, $i=1,...,q$, and $\sigma^{2}_{\epsilon}$ by $\tilde{\beta}_{i}$, $i=1,...,q$, and $\tilde{\sigma}^{init}_{\epsilon}$ respectively. We replace $\sigma^{2}_{\eta}$ with its consistent point estimate $\hat{\sigma}^{2}_{\eta}$. Thus, the updated fiducial pivots for $\alpha_{i}$, $i=1,...,q$, can be obtained as
$\tilde{\alpha}_{i}=\frac{\sum\limits_{j=2}^{r}\sum\limits_{k=1}^{N_{ij}}\frac{y_{ijk}}{x_{j}}-\tilde{\beta}_{i}\sum\limits_{j=2}^{r}\sum\limits_{k=1}^{N_{ij}}\exp(\hat{\sigma}_{\eta}z_{ijk}^{\eta})-\tilde{\sigma}^{init}_{\epsilon}\sum\limits_{j=2}^{r}\sum\limits_{k=1}^{N_{ij}}\frac{z_{ijk}^{\epsilon}}{x_{j}}}{N_{ij}\sum\limits_{j=2}^{r}\frac{1}{x_{j}}},\quad i=1,...,q.$ (5)
4.1.4 Fiducial Quantity for $\sigma^{2}_{\eta}$ and Updated Fiducial Quantity $\sigma^{2}_{\epsilon}$
@@
$\sum\limits_{i=1}^{q}\sum\limits_{k=1}^{N_{i1}}\left(y_{i1k}-\frac{1}{q}\sum\limits_{i=1}^{q}\frac{1}{N_{ij}}\sum\limits_{k=1}^{N_{i1}}y_{i1k}\right)^{2}+\sum\limits_{j=2}^{r}\sum\limits_{i=1}^{q}\sum\limits_{k=1}^{N_{ij}}\left(\frac{y_{ijk}}{x_{j}}-\frac{1}{q}\sum\limits_{i=1}^{q}\frac{1}{N_{ij}}\sum\limits_{k=1}^{N_{ij}}\frac{y_{ijk}}{x_{j}}\right)^{2}=\\ \sum\limits_{i=1}^{q}\sum\limits_{k=1}^{N_{i1}}\left(\tilde{\alpha}_{i}+\sigma_{\epsilon}z_{i1k}^{\epsilon}-\frac{1}{q}\sum\limits_{i=1}^{q}\frac{1}{N_{i1}}\sum\limits_{k=1}^{N_{i1}}\left(\tilde{\alpha}_{i}+\sigma_{\epsilon}z_{i1k}^{\epsilon}\right)\right)^{2}+\sum\limits_{j=2}^{r}\sum\limits_{i=1}^{q}\sum\limits_{k=1}^{N_{ij}}\\ \left(\frac{\tilde{\alpha}_{i}}{x_{j}}+\tilde{\beta}_{i}exp(\sigma_{\eta}z_{ijk}^{\eta})+\sigma_{\epsilon}\frac{z_{ijk}^{\epsilon}}{x_{j}}-\frac{1}{q}\sum\limits_{i=1}^{q}\frac{1}{N_{ij}}\sum\limits_{k=1}^{N_{ij}}\left(\frac{\tilde{\alpha}_{i}}{x_{j}}+\tilde{\beta}_{i}exp(\sigma_{\eta}z_{ijk}^{\eta})+\sigma_{\epsilon}\frac{z_{ijk}^{\epsilon}}{x_{j}}\right)\right)^{2}.$ (7)
In the above structural equations, $\alpha_{i}$, $\beta_{i}$, $i=1,...,q$, are replaced by their corresponding fiducial quantities $\tilde{\alpha}_{i}$, $\tilde{\beta}_{i}$, $i=1,...,q$, respectively. These structural equations are nonlinear in $\sigma_{\eta}^{2}$ and $\sigma_{\epsilon}^{2}$, and can be solved numerically to obtain a fiducial quantity for $\sigma_{\eta}^{2}$, and an updated fiducial quantity for $\sigma_{\epsilon}^{2}$; denoted by $\tilde{\sigma}_{\eta}^{2}$ and $\tilde{\sigma}_{\epsilon}^{2}$, respectively.
4.2 Fiducial Quantity for an Unknown Concentration
@@
Algorithm 1 Generating Fiducial Pivot Quantities for Unknown Concentrations
@@
2:$N$ fiducial pivot quantities for unknown concentrations $\mathcal{X}_{j}$, $j=1,...,r^{*}$.
@@
4:     Generate $z^{\eta}_{ijk}\overset{i.i.d.}{\sim}N(0,1)$, $z^{\epsilon}_{ijk}\overset{i.i.d.}{\sim}N(0,1)$, for $i=1,...,q$, $j=1,...,r$, $k=1,...,N_{ij}$.
@@
7:     Generate fiducial quantity for $\sigma_{\eta}$ and updated fiducial quantity for $\sigma_{\epsilon}$ by solving equations (6) and (7) simultaneously.
@@
9:     Obtain fiducial quantity of unknown concentration $\mathcal{X}_{j}$ using expression (9) for each $j=1,...,r^{*}$.
@@
11:return $N$ fiducial pivot quantities for each unknown concentration $\mathcal{X}_{j}$, $j=1,...,r^{*}$ from step 7.
@@ 4.3 Fiducial Confidence Interval and Fiducial Point Estimate
Since fiducial quantities are functions of the observed data and random quantities, whose distributions are free of any unknown parameters, samples from the fiducial quantities can be obtained simply by drawing samples from the corresponding distributions. These samples can then be used to estimate the density of the fiducial distribution for each parameter. Once the density is estimated, it serves as a basis for determining both point and interval estimates for each parameter. By examining the estimated density, we can derive point estimates such as the mode, and the highest density intervals, which offer a measure of uncertainty around these estimates. In this work, we focus on the point and interval estimation of the unknown concentration from new samples $\{y^{*}_{ijk}:i=1,...,q^{*},j=1,...,r^{*},k=1,...,N_{ij}^{*}\}$, where the fiducial quantities of the model parameters have been obtained from the earlier sample $\{y_{ijk}:i=1,...,q,j=1,...,r,k=1,...,N_{ij}\}$, corresponding to known concentrations. For an unknown concentration $\mathcal{X}_{j}$ (for any $j=1,..,r^{*}$), one sample from its fiducial distribution can be drawn sequentially based on the discussion in sections 4.1 and 4.2 using algorithm 1. One can generate $N$ fiducial quantities of each unknown concentration by replicating these steps $N$ times. Let us denote the density of the fiducial distribution of the unknown concentration $\mathcal{X}_{j}$ as $d_{\mathcal{X}_{j}}(.)$ which can be estimated accurately from the $N$ generated fiducial quantities. In the following, we use the estimated 95% fiducial highest density interval as the interval estimate and the fiducial mode as the point estimate of $\mathcal{X}_{j}$.
100$(1-\alpha)$Fiducial Highest Density Interval (HDI). The 100$(1-\alpha)$fiducial HDI for $\mathcal{X}_{j}$ is the interval $[L,U]$ such that: $\int\limits_{L}^{U}d_{\mathcal{X}_{j}}(x)\delta x=1-\alpha$ and minimizes the interval length $(U-L)$. The endpoints $L$ and $U$ are determined such that $d_{\mathcal{X}_{j}}(L)=d_{\mathcal{X}_{j}}(U)$ ensuring that the interval captures the highest density region. Fiducial Mode. The fiducial mode $\hat{\mathcal{X}_{j}}$ is the value of $\mathcal{X}_{j}$ that maximizes the fiducial density function; i.e., $\hat{\mathcal{X}_{j}}=\arg\max_{x}d_{\mathcal{X}_{j}}(x)$. This point estimate $\hat{\mathcal{X}_{j}}$ represents the value of $\mathcal{X}_{j}$ with the highest probability density under the fiducial distribution.
@@ 5.2 Simulation Results
Interval Estimation: The simulation results for each scenario are based on 5,000 simulated datasets. For each dataset, the proposed fiducial method estimates the unknown concentration using 10,000 samples generated from the fiducial quantity. Similarly, 10,000 bootstrap samples are utilized in both the parametric bootstrap and LRT-based methods for comparison. Notably, the LRT-based method is computationally more intensive than the parametric bootstrap method. However, the proposed fiducial method provides significantly more computational efficiency compared to both approaches. The simulation results for interval estimation are presented in Tables 1, 2, 3, 4. For simulation scenarios 1.B and 2.B (Tables 2 and 4), results for concentrations 0, 15, 30, and 50 are presented in this document (Complete tables can be found in Supplementary Material). The performance of the interval estimate is evaluated based on both the average width of the 95% confidence interval and its coverage probability.
Figure 1: Density plots of the fiducial distribution, bootstrap distribution and Wald-Based normal distribution for an unknown concentration (true value is 50) for a sample from simulation setting 2.A where the MLE is 55.6. Vertical dashed lines shows 95% confidence intervals. The figure highlights the fact that only the fiducial interval includes the true value: a conclusion that agrees with the poor coverage probabilities of the other intervals, as reported in this section.
@@
(×$\times 10^{2}$) (×$\times 10^{2}$) (×$\times 10^{2}$) (×$\times 10^{2}$) (×$\times 10^{2}$) (×$\times 10^{2}$)
@@
(×$\times 10^{2}$) (×$\times 10^{2}$) (×$\times 10^{2}$) (×$\times 10^{2}$) (×$\times 10^{2}$) (×$\times 10^{2}$)
@@ 5.3 Computational Efficiency of the Fiducial Approach
Computational efficiency should be an important factor when developing new methods, especially for interdisciplinary applications, as these methods may be integrated into publicly available software and will be adopted by researchers across disciplines. Thus, in addition to demonstrating the superiority of the fiducial confidence intervals, we would like to highlight its computational advantage over the other available methods. The fiducial approach entails drawing fiducial samples; for this, we have to numerically solve only two equations (for $\sigma_{\eta}$ and $\sigma_{\epsilon}$) since closed-form expressions are available for the fiducial pivotal quantities for all other model parameters ($\alpha_{i},\beta_{i},i=1,...,q$), as well as the unknown concentrations $\mathcal{X}_{j},j=1,...,r$. Thus, obtaining a fiducial sample is computationally very cheap. All other approaches, including the bootstrap, require extensive maximum likelihood calculations, including numerical integration and numerical optimization of the integrated likelihood. This can be computationally extensive compared to the proposed fiducial method, especially when the data come from multiple laboratories. The computation is especially burdensome for the modified likelihood ratio method due to Zhao et al. (2021) since it involves likelihood maximization and bootstrap sampling at several stages.
@@
Let us denote the computation time for the bootstrap method as ‘$t$’. Now, coming into Zhao’s LRT, it needs a large number of bootstrap samples and computation of constrained and unconstrained MLEs to obtain a single value of the standardized LRT statistics. Note that this computational time is ‘$2t$’ (each ‘$t$’ for constrained and unconstrained MLE). Further, obtaining each limit of the confidence interval requires the generation of ‘$k$’ values of the standardized LRT. Thus, a confidence interval of a single unknown concentration requires computation time ‘$4kt$’. To achieve decent accuracy, a value of ‘$k$’ is considered at least 50. Additionally, these samples can not be used again to obtain intervals for another unknown concentration, even if the samples come from the same lab. Now, if there are 10 unknown concentrations of interest, the computation time will increase simply 10 times. The proposed fiducial method can utilize the same fiducial pivot of the model parameters to obtain pivots for all unknown concentrations. In summary, to obtain intervals for 10 unknown concentrations, if the computation time for bootstrap is ‘$t$’, LRT takes ‘$2000t$’, whereas fiducial samples are take less than ‘$t$’, as a fiducial sample is computationally much cheaper than a bootstrap sample. The above discussion shows the computational extensiveness of Zhao’s method. It becomes computationally impractical to evaluate its performance in simulation cases with multiple laboratories and many unknown concentrations. This is one of the reasons why it is not being used in simulation cases apart from simulation scenario 1.A.

### v3

@@ 4.2.3 Updated Fiducial Quantities for $\alpha_{i}$
The method for constructing equations that lead to fiducial quantities for $\alpha_{i}$ is similar to that for $\beta_{i}$ (for i=1,,qi=1,...,q), as discussed in the previous subsection. In the following, we construct a total of $q$ equations, one for each $\alpha_{i}$, by summing equation (4) over all replications at different non-zero concentrations for each lab, $\sum\limits_{j=2}^{r}\sum\limits_{k=1}^{N_{ij}}\frac{y_{ijk}}{x_{j}}=\alpha_{i}\sum\limits_{j=2}^{r}N_{ij}\frac{1}{x_{j}}+\beta_{i}\sum\limits_{j=2}^{r}\sum\limits_{k=1}^{N_{ij}}\exp(\sigma_{\eta}z_{ijk}^{\eta})+\sigma_{\epsilon}\sum\limits_{j=2}^{r}\sum\limits_{k=1}^{N_{ij}}\frac{z_{ijk}^{\epsilon}}{x_{j}},\ i=1,...,q.$ Again, these are linear in each $\alpha_{i}$, i=1,,qi=1,...,q. We now replace $\beta_{i}$, i=1,,qi=1,...,q, and $\sigma^{2}_{\epsilon}$ by $\tilde{\beta}_{i}$, i=1,,qi=1,...,q, and $\tilde{\sigma}^{init}_{\epsilon}$ respectively. We replace $\sigma^{2}_{\eta}$ with its consistent point estimate $\hat{\sigma}^{2}_{\eta}$. Thus, the updated fiducial pivotal quantities for $\alpha_{i}$, i=1,,qi=1,...,q, can be obtained as
$\tilde{\alpha}_{i}=\frac{\sum\limits_{j=2}^{r}\sum\limits_{k=1}^{N_{ij}}\frac{y_{ijk}}{x_{j}}-\tilde{\beta}_{i}\sum\limits_{j=2}^{r}\sum\limits_{k=1}^{N_{ij}}\exp(\hat{\sigma}_{\eta}z_{ijk}^{\eta})-\tilde{\sigma}^{init}_{\epsilon}\sum\limits_{j=2}^{r}\sum\limits_{k=1}^{N_{ij}}\frac{z_{ijk}^{\epsilon}}{x_{j}}}{\sum\limits_{j=2}^{r}N_{ij}\frac{1}{x_{j}}},\quad i=1,...,q.$ (5)
4.2.4 Fiducial Pivotal Quantity for $\sigma^{2}_{\eta}$ and Updated Fiducial Pivotal Quantity $\sigma^{2}_{\epsilon}$
@@
$\sum\limits_{i=1}^{q}\sum\limits_{k=1}^{N_{i1}}\left(y_{i1k}-\frac{1}{q}\sum\limits_{i=1}^{q}\frac{1}{N_{i1}}\sum\limits_{k=1}^{N_{i1}}y_{i1k}\right)^{2}+\sum\limits_{j=2}^{r}\sum\limits_{i=1}^{q}\sum\limits_{k=1}^{N_{ij}}\left(\frac{y_{ijk}}{x_{j}}-\frac{1}{q}\sum\limits_{i=1}^{q}\frac{1}{N_{ij}}\sum\limits_{k=1}^{N_{ij}}\frac{y_{ijk}}{x_{j}}\right)^{2}=\\ \sum\limits_{i=1}^{q}\sum\limits_{k=1}^{N_{i1}}\left(\tilde{\alpha}_{i}+\sigma_{\epsilon}z_{i1k}^{\epsilon}-\frac{1}{q}\sum\limits_{i=1}^{q}\frac{1}{N_{i1}}\sum\limits_{k=1}^{N_{i1}}\left(\tilde{\alpha}_{i}+\sigma_{\epsilon}z_{i1k}^{\epsilon}\right)\right)^{2}+\sum\limits_{j=2}^{r}\sum\limits_{i=1}^{q}\sum\limits_{k=1}^{N_{ij}}\\ \left(\frac{\tilde{\alpha}_{i}}{x_{j}}+\tilde{\beta}_{i}exp(\sigma_{\eta}z_{ijk}^{\eta})+\sigma_{\epsilon}\frac{z_{ijk}^{\epsilon}}{x_{j}}-\frac{1}{q}\sum\limits_{i=1}^{q}\frac{1}{N_{ij}}\sum\limits_{k=1}^{N_{ij}}\left(\frac{\tilde{\alpha}_{i}}{x_{j}}+\tilde{\beta}_{i}exp(\sigma_{\eta}z_{ijk}^{\eta})+\sigma_{\epsilon}\frac{z_{ijk}^{\epsilon}}{x_{j}}\right)\right)^{2}.$ (7)
In the above structural equations, $\alpha_{i}$, $\beta_{i}$, i=1,,qi=1,...,q, are replaced by their corresponding fiducial quantities $\tilde{\alpha}_{i}$, $\tilde{\beta}_{i}$, i=1,,qi=1,...,q, respectively. These structural equations are nonlinear in $\sigma_{\eta}^{2}$ and $\sigma_{\epsilon}^{2}$, and can be solved numerically to obtain a fiducial pivotal quantity for $\sigma_{\eta}^{2}$, and an updated fiducial pivotal quantity for $\sigma_{\epsilon}^{2}$; denoted by $\tilde{\sigma}_{\eta}^{2}$ and $\tilde{\sigma}_{\epsilon}^{2}$, respectively.
Numerical solutions: Existence and uniqueness.
For each fiducial draw, equations (6)-(7) define a two-dimensional nonlinear system in $(\sigma_{\eta},\sigma_{\epsilon})$ whose properties depend on the realized auxiliary variables $\bm{z}^{\eta}=\{z^{\eta}_{ijk}\}$ and $\bm{z}^{\epsilon}=\{z^{\epsilon}_{ijk}\}$ (with the observed data $\bm{y}=\{y_{ijk}\}$ treated as fixed). Writing the system as $E_{1}(\sigma_{\eta},\sigma_{\epsilon},\bm{z}^{\eta},\bm{z}^{\epsilon};\bm{y})=0$ in (6) and $E_{2}(\sigma_{\eta},\sigma_{\epsilon},\bm{z}^{\eta},\bm{z}^{\epsilon};\bm{y})=0$ in (7), we note that global existence and global uniqueness are generally difficult to guarantee for coupled nonlinear equations. However, a standard sufficient condition for local uniqueness can be stated: if a solution exists and the Jacobian matrix $J(\sigma_{\eta},\sigma_{\epsilon})=\partial(E_{1},E_{2})/\partial(\sigma_{\eta},\sigma_{\epsilon})$ is nonsingular at that solution (i.e., $\det\{J(\sigma_{\eta},\sigma_{\epsilon})\}\neq 0$), then the solution is locally unique (by the implicit function theorem). Whether this condition holds depends on $(\bm{z}^{\eta},\bm{z}^{\epsilon})$ and the observed data.
In practice, since (6)-(7) involves only two parameters, we solve the system numerically using a standard root-finding routine with multiple starting values (a small $3\times 3$ grid). If a draw does not yield a solution from any starting value, we do not include that draw in the fiducial sample used for inference. If multiple solutions are returned, we select the solution that minimizes the residual sum of squares $\sum\limits_{i,j,k}\{y_{ijk}-(\tilde{\alpha}_{i}+\tilde{\beta}_{i}x_{j}\exp(\sigma_{\eta}z^{\eta}_{ijk})+\sigma_{\epsilon}z^{\epsilon}_{ijk})\}^{2}$. Empirically, we have observed that draws with no solution or multiple solutions become less frequent as the total sample size $\sum\limits_{i,j}N_{ij}$ increases (e.g., under the larger-sample simulation configurations such as Scenarios 1.B and 2.B).
4.3 Fiducial Pivotal Quantity for an Unknown Concentration
@@
Algorithm 1 Generating Fiducial Pivotal Quantities for Unknown Concentrations
@@
2:NN fiducial pivotal quantities for unknown concentrations $\mathcal{X}_{j}$, $j=1,...,r^{*}$.
@@
4:  Generate $z^{\eta}_{ijk}\overset{i.i.d.}{\sim}N(0,1)$, $z^{\epsilon}_{ijk}\overset{i.i.d.}{\sim}N(0,1)$, for i=1,,qi=1,...,q, j=1,,rj=1,...,r, $k=1,...,N_{ij}$.
@@
7:  Generate fiducial pivotal quantity for $\sigma_{\eta}$ and updated fiducial pivotal quantity for $\sigma_{\epsilon}$ by solving equations (6) and (7) simultaneously.
@@
9:  Obtain fiducial pivotal quantity of unknown concentration $\mathcal{X}_{j}$ using expression (9) for each $j=1,...,r^{*}$.
@@
11:return NN fiducial pivotal quantities for each unknown concentration $\mathcal{X}_{j}$, $j=1,...,r^{*}$ from step 7.
@@ 4.4 Fiducial Confidence Interval and Fiducial Point Estimate
Since fiducial quantities are functions of the observed data and random quantities, whose distributions are free of any unknown parameters, samples from the fiducial quantities can be obtained simply by drawing samples from the corresponding distributions. These samples can then be used to estimate the density of the fiducial distribution for each parameter. Once the density is estimated, it serves as a basis for determining both point and interval estimates for each parameter. By examining the estimated density, we can derive point estimates such as the mode, and the highest density intervals, which offer a measure of uncertainty around these estimates. In this work, we focus on the point and interval estimation of the unknown concentration from new samples $\{y^{*}_{ijk}:i=1,...,q^{*},j=1,...,r^{*},k=1,...,N_{ij}^{*}\}$, where the fiducial quantities of the model parameters have been obtained from the earlier sample $\{y_{ijk}:i=1,...,q,j=1,...,r,k=1,...,N_{ij}\}$, corresponding to known concentrations. For an unknown concentration $\mathcal{X}_{j}$ (for any $j=1,..,r^{*}$), one sample from its fiducial distribution can be drawn sequentially based on the discussion in sections 4.2 and 4.3 using algorithm 1. One can generate $N$ fiducial quantities of each unknown concentration by replicating these steps $N$ times. Let us denote the density of the fiducial distribution of the unknown concentration $\mathcal{X}_{j}$ as $d_{\mathcal{X}_{j}}(.)$ which can be estimated accurately from the $N$ generated fiducial quantities. In the following, we use the estimated 95% fiducial highest density interval as the interval estimate and the fiducial mode as the point estimate of $\mathcal{X}_{j}$.
[…]
Fiducial Mode. The fiducial mode $\hat{\mathcal{X}_{j}}$ is the value of $\mathcal{X}_{j}$ that maximizes the fiducial density function; i.e., $\hat{\mathcal{X}_{j}}=\arg\max_{x}d_{\mathcal{X}_{j}}(x)$. This point estimate $\hat{\mathcal{X}_{j}}$ represents the value of $\mathcal{X}_{j}$ with the highest probability density under the fiducial distribution.
@@ 5.2 Simulation Results
Interval Estimation: The simulation results for each scenario are based on 5,000 simulated datasets. For each dataset, the proposed fiducial method estimates the unknown concentration using 10,000 samples generated from the fiducial pivotal quantity. Similarly, 10,000 bootstrap samples are utilized in both the parametric bootstrap and LRT-based methods for comparison. Notably, the LRT-based method is computationally more intensive than the parametric bootstrap method. However, the proposed fiducial method provides significantly more computational efficiency compared to both approaches. The simulation results for interval estimation are presented in Tables 1, 2, 3, 4. For simulation scenarios 1.B and 2.B (Tables 2 and 4), results for concentrations 0, 15, 30, and 50 are presented in this document (Complete tables can be found in Supplementary Material). The performance of the interval estimate is evaluated based on both the average width of the 95% confidence interval and its coverage probability.
Figure 1: Density plots of the fiducial distribution, bootstrap distribution and Wald-Based normal distribution for an unknown concentration (true value is 50) for a sample from simulation setting 2.A where the MLE is 55.6. Vertical dashed lines shows 95% confidence intervals. The figure highlights the fact that only the fiducial interval includes the true value: a conclusion that agrees with the poor coverage probabilities of the other intervals, as reported in this section. Note that in the overlaid contour/density display, apparent colour changes at intersections can occur as a rendering artifact of overlapping densities and do not represent an additional method.
@@
($\times 10^{2}$) ($\times 10^{2}$) ($\times 10^{2}$) ($\times 10^{2}$) ($\times 10^{2}$) ($\times 10^{2}$)
@@
($\times 10^{2}$) ($\times 10^{2}$) ($\times 10^{2}$) ($\times 10^{2}$) ($\times 10^{2}$) ($\times 10^{2}$)
@@ 5.3 Computational Efficiency of the Fiducial Approach
Computational efficiency should be an important factor when developing new methods, especially for interdisciplinary applications, as these methods may be integrated into publicly available software and will be adopted by researchers across disciplines. Thus, in addition to demonstrating the superiority of the fiducial confidence intervals, we highlight its computational advantage over the other available methods. The fiducial approach entails drawing fiducial samples; for this, we have to numerically solve only two equations (for $\sigma_{\eta}$ and $\sigma_{\epsilon}$) since closed-form expressions are available for the fiducial pivotal quantities for all other model parameters ($\alpha_{i},\beta_{i},i=1,...,q$), as well as the unknown concentrations $\mathcal{X}_{j},j=1,...,r$. Thus, obtaining a fiducial sample is computationally very cheap. All other approaches, including the bootstrap, require extensive maximum likelihood calculations, including numerical integration and numerical optimization of the integrated likelihood. This can be computationally extensive compared to the proposed fiducial method, especially when the data come from multiple laboratories. The computation is especially burdensome for the modified likelihood ratio method due to Zhao et al. (2021) since it involves likelihood maximization and bootstrap sampling at several stages.
@@
Let us denote the computation time for the bootstrap method as ‘$t$’. Now, coming into Zhao’s LRT, it needs a large number of bootstrap samples and computation of constrained and unconstrained MLEs to obtain a single value of the standardized LRT statistics. Note that this computational time is ‘$2t$’ (each ‘$t$’ for constrained and unconstrained MLE). Further, obtaining each limit of the confidence interval requires the generation of ‘$k$’ values of the standardized LRT. Thus, a confidence interval of a single unknown concentration requires computation time ‘$4kt$’. To achieve decent accuracy, a value of ‘$k$’ is considered at least 50. Additionally, these samples can not be used again to obtain intervals for another unknown concentration, even if the samples come from the same lab. Now, if there are 10 unknown concentrations of interest, the computation time will increase simply 10 times. The proposed fiducial method can utilize the same fiducial pivotal quantity of the model parameters to obtain fiducial pivotal quantities for all unknown concentrations. In summary, to obtain intervals for 10 unknown concentrations, if the computation time for bootstrap is ‘$t$’, LRT takes ‘$2000t$’, whereas fiducial samples are take less than ‘$t$’, as a fiducial sample is computationally much cheaper than a bootstrap sample. The above discussion shows the computational extensiveness of Zhao’s method. It becomes computationally impractical to evaluate its performance in simulation cases with multiple laboratories and many unknown concentrations. This is one of the reasons why it is not being used in simulation cases apart from simulation scenario 1.A.
Example of wall-clock runtimes.
To complement the computational comparison in terms of $t$, we also report wall-clock runtimes. All computations were carried out on a laptop with Intel(R) Core(TM) i7-10510U CPU @ 1.80GHz (2.30GHz) and 16GB RAM. Using 1000 resamples/samples, the average bootstrap runtime is approximately $t\approx 320$ seconds in the smaller-sample scenarios (1.A and 2.A), and $t\approx 3680$ seconds in the larger-sample scenarios (1.B and 2.B). In contrast, the proposed fiducial method requires approximately 241 seconds on average in the larger-sample scenarios using 1000 fiducial samples. Translating the scaling argument above, for 10 unknown concentrations the modified LRT method of Zhao et al. (2021) would require approximately $2000t$, which corresponds to about $7.36\times 10^{6}$ seconds (approximately 85 days) when $t=3680$ seconds in the larger-sample setting, illustrating its computational impracticality for the multi-laboratory/multiple-unknown regime studied here.